OpenStudy (anonymous):

5x + y = 6 5x + 3y = -4 The y-coordinate of the solution to the system shown is _____. -5 -1 1/2 got -1

OpenStudy (anonymous):

Did you solve the top equation for y? y=6-5x?

OpenStudy (anonymous):

Then substitute that into the second equation: 5x + 3(6-5x)=-4

OpenStudy (campbell_st):

-1 is not correct.... just subtract the 2nd equation from the 1st equation to eliminate x 5x + y = 6 - 5x + 3y = -4 ---------------- -2y = 10 y = -5 work out which pronumeral needs to be eliminated before starting the question... and work to that